### v1

@@ 2.6 Reward Modeling Approaches
For each epoch $e_{i}$ and the fixed price path $\bm{P}_{e_{i}}$, random $\tau$–symmetric strategies are drawn from $\mathbf{S}_{e_{i}}^{\tau}$. The optimal strategy $\varphi_{e_{i}}^{\mathrm{OPT}}$ maximizes the income from the LP’s fee and, by construction, coincides with the pool model fee: $\bm{L}_{e_{i}}^{\Sigma}$ is therefore concentrated in $12\tau+$ 1 ranges, mirroring Stage 2 from Section 2.4 but at the pool level rather than for an individual LP. Figure 6 illustrates how the liquidity state in each bucket changes while historical prices evolve under one realization of the $\varphi^{\tau}$ strategy.
@@
With the initial LP capital $W_{e_{i}}^{\mathrm{LP}}$ and the same price subset $\bm{P}_{e_{i}}$, a random search of the strategy family $\mathbf{S}_{e_{i}}^{\tau}$ (for the fixed parameter $\tau$) produces the optimal strategy $\varphi_{e_{i}}^{\mathrm{OPT}}$ that maximizes the LP reward $f_{e_{i}}^{\mathrm{LP}}$ within the liquidity of the approximated pool, cf.(22). Figure 8 illustrates how the liquidity state in each bucket changes while historical prices evolve, and $\bm{L}_{e_{i}}^{\Sigma}$ in the $12\tau+$ 1 buckets of $\beta$ contains, among other components, the LP-position in the form of $\bm{L}_{e_{i}}^{\mathrm{LP}}$ dictated by $\varphi_{e_{i}}^{\mathrm{OPT}}$.
@@
The distinguishing element of the present approach is that LP fees are calculated after local expansion of historical liquidity by the LP’s own liquidity $\bm{L}_{e_{i}}^{\mathrm{LP}}$ into $12\tau+$ 1 buckets centered on the reference index $M$. Thus, the fee share is taken relative to the augmented liquidity profile - historical plus the additional capital of the LP - see Fig. 10.
@@
3 Related Work
This section includes a number of papers studying liquidity provision in CLMM markets, specific strategies and backtest frameworks, as well as papers on ML applications for CLMM markets.
The concept of liquidity provision in the context of concentrated liquidity has been studied in the literature from several angles. Popular research topics are impermanent loss, trading fees in the DEX market, DEX market overall design, liquidity provision strategies, and many others.
The liquidity provided into a DEX pool both for the CPMM and CLMM cases is an option-like property with the corresponding nonlinear payoff function and effects caused by volatility and elapsed time. The general results behind an abstract liquidity pool have been developed in [7], and [9]. There, one can find all the necessary calculations to deal with token reserves, swaps, fees, and liquidity. Moreover, in [10] the authors provide a methodology to estimate the accumulated fees in a liquidity pool with basic assumptions on the price behavior, but do not explicitly consider the trading volumes.
To understand the properties of a decentralized exchange from the point of view of a trader, consider the paper [11] in which the authors compare the price discovery process in a DEX with the centralized market and study the behavior of informed traders. In the paper [12] the authors work with the bid and ask prices in the context of a liquidity pool and derive the value of a no-trade gap.
Among the important papers which study the impermanent loss one should mention the paper [13] where the authors study the risks of a liquidity provider and the potential returns which may be obtained from such activity. The model allows to dive into the liquidity provision problem from the point of view of a cautious investor and allows to look at the numerical estimations from the real-world data developing solid intuition to deal with such kind of problems.
The next paper to consider is [14] where the authors provide an estimate that half of the liquidity providers lose money in the dataset considered in the paper. This is a reminder that liquidity provision is very complicated in the case of concentrated liquidity, and one must deal with option like payoffs as well as with dynamic reallocation decisions to capture the fees depending on the market conditions.
To understand the concept of impermanent loss even deeper one should dive into the [15] paper which not only provides a concise and instructive result to estimate the impermanent loss but also generalizes the concept allowing to compare the potential loss not only to the buy and hold strategy but also to any benchmark strategy available to the investor with a special case of the rebalancing strategy producing the corresponding loss versus rebalancing metric.
Dealing with impermanent loss is a complicated problem, a general approach is hedging the risk with classic options. The following study [16] applies arbitrage-based methods to produce both a static and a dynamic hedge for the impermanent loss and derive explicit formulae in the case of a Black-Scholes-Merton model and the log-normal stochastic volatility model.
The concepts described above are the important building blocks necessary to construct a liquidity provision strategy for a CLMM market. The strategies studied in the paper are based on the concept of $\tau$-reset described in [5].
A notable paper studying liquidity provision strategies is [17] where the authors consider a CLMM and consider the accumulated fees, rebalancing costs and the liquidity concentration risk. This is a continuous time model which allows to analyze asymmetric liquidity provision strategies for specific market conditions.
The paper [18] considers the problems arising in the liquidity pools in continuous time. The authors consider a continuous time model for the liquidity dynamics in a CLMM pool. Both finite and infinite horizon settings are considered with an assumption that the liquidity provider should rebalance the strategy when the price goes out of the price range. Following this research it is also important to mention a very promising piece of research considering liquidity provision as an optimal stopping problem [19]. This approach allows to look at the liquidity provision problem from a new point of view and is important to consider for the general idea of a dynamic liquidity provision strategy.
The papers described above focus on modeling and developing theoretical results meanwhile an efficient way to work with liquidity provision strategies is based on numerical analysis and machine learning techniques. Several papers study CLMM liquidity provision from this point of view, for example in the paper [20] the authors apply the deep reinforcement learning approach to maximize the accumulated fees corrected for the loss versus rebalancing risk, and optimize the price interval width, lower and upper bound of the interval. Another paper with a similar strategy and deep reinforcement learning approach is studied in [21], where the authors demonstrate better strategy performance compared to a benchmark on the example of ETH/USDC and ETH/USDT pools.
The reinforcement learning approach is applied not only to liquidity provision strategies but for the AMM design to create a more efficient market decreasing the impermanent loss and slippage, as shown in [22].
Efficient decentralized market design is another important topic to understand the ideas behind optimal liquidity provision. The paper [23] introduces the concept of a general decentralized liquidity pool and allows the liquidity provider to use dynamic bonding curves for more efficient trading. And the paper [24] extends the problem even further considering the trading platform itself as another player competing for the trading fees with the liquidity providers.
In the context of current research, it is important to remember that the liquidity provider is not a single player in the pool, there are many different liquidity providers who compete to earn more fees while controlling for the risk of impermanent loss. In the paper [25] the authors apply a mean-field game approach to analyze the resulting liquidity distribution which appears as the result of many liquidity provision strategies and also consider the case of just-in-time liquidity as an extension of the model. A model specifically focused on just-in-time liquidity is studied in [26].
The research papers discussed above allow to achieve deep understanding of the liquidity provision problem, the only element left out in the context of strategies is backtesting. To understand whether the strategy is of any use for practice one must launch it in a real liquidity pool or consider a robust simulation environment allowing to estimate the amount of fees accumulated by the strategy in the pool. General ideas behind the problem of strategy backtesting for CLMMs have been studied thoroughly in our previous paper, we refer the reader to the paper [6] for deeper analysis.
In this paper, we apply the accumulated knowledge of the concentrated liquidity pools mechanics to extend the backtesting framework research and use it for better liquidity provision activity using machine learning methods for to obtain numerical results for the strategies in specific market conditions within the family of $\tau$-reset strategies.
@@ 4 Modeling on Real Data
In this section, the methodology developed above is applied to real Uniswap V3 data. A uniform-liquidity strategy is used as a benchmark, while the proposed approach is evaluated on data from USDC/ETH pools with several fee levels, in stablecoin pools and in WBTC/ETH. The section ends with (i) a summary of the consolidated performance of all pools, (ii) a comparison with an external backtesting tool, and (iii) a high-level schematic of the application of the ML model for liquidity allocation decisions.
@@
Historical swap-transaction data were retrieved using GraphQL queries to the Uniswap V3 subgraph.
@@ 4.1 Historical Liquidity Approximation and Optimal LP Strategy Search
[…]
@@
[…]
@@ 4.2 ML Model Training and Features Engineering
Feature engineering is also intentionally minimal. All inputs are derived from only two raw numerical indicators, from which a compact set of features is generated to capture latent patterns in market dynamics. The approach is generic and can be adapted to any collection of numeric variables. Subsequent sections confirm that even such sparse feature sets and simple model configurations are sufficient for the goals of this work.
@@
The predictions of the integral model are produced by a weighted ensemble of three base learners: an MLP neural network, CatBoost, and an LSTM network. The remainder of this subsection details the architectures implemented in the experiments, including structural choices, theoretical rationale, hyperparameter settings, regularization techniques, and evaluation metrics.
@@
  • Output layer: Dense$(\tau+1,\ \text{activation}=\texttt{softmax})$
@@
  • Output layer: Dense$(\tau+1,\ \text{activation}=\texttt{softmax})$
@@
The final prediction is obtained by a weighted sum of the three base models. The optimal weight vector $\bm{w}=\{w_{1},w_{2},w_{3}\}$ is found by minimizing the mean square error with regularization $L_{2}$ ($\alpha=0.001$) under the constraints $w_{j}>0$ and $\sum_{j=1}^{3}w_{j}=1$.
@@ 4.3 Feature Engineering
Centralized exchange (CEX) data serve as the information set to characterize the market state in each allocation decision. Historical market data are sourced through the live_trading_indicators [^2] https://pypi.org/project/live-trading-indicators/ library, which offers unified access to multiple cryptocurrency exchanges. For the USDC/ETH pool, the study employs one-minute OHLCV bars for the USDC/USDT and BTC/USDT pairs on Binance, covering 1 March 2023 to 30 September 2024; the inclusion of BTC captures potential cross-asset comovement with ETH. For other pools, the same procedure is followed: features are generated for both the target pair and the BTC.
@@ 4.4 Application of ML Models to the OOT Sample
Figure 17: Accumulated LP fees depending on $\tau$ (USDC/ETH 0.3%).
@@
[…]
@@
Throughout the OOT backtesting, all fees are not reinvested between epochs; this separation isolates the modeled reward stream from the evolving principal $W^{\mathrm{LP}}$.
@@
(first pool USDC/ETH 0.3%)
@@ 4.5 Comparison with an Alternative Tool
Range Size 10 USDC 10 USDC 1.23e-4 WBTC 1.57e-4 USDC

### v2

@@ 4.3 Reward Modeling Approaches
For each epoch $e_{i}$ and the fixed price path $\bm{P}_{e_{i}}$, random $\tau$–symmetric strategies are drawn from $\mathbf{S}_{e_{i}}^{\tau}$. The optimal strategy $\varphi_{e_{i}}^{\mathrm{OPT}}$ maximizes the income from the LP’s fee and, by construction, coincides with the pool model fee: $\bm{L}_{e_{i}}^{\Sigma}$ is therefore concentrated in $2\tau+1$ ranges, mirroring Stage 2 from Section 4.1 but at the pool level rather than for an individual LP. Figure 6 illustrates how the liquidity state in each bucket changes while historical prices evolve under one realization of the $\varphi^{\tau}$ strategy.
@@
With the initial LP capital $W_{e_{i}}^{\mathrm{LP}}$ and the same price subset $\bm{P}_{e_{i}}$, a random search of the strategy family $\mathbf{S}_{e_{i}}^{\tau}$ (for the fixed parameter $\tau$) produces the optimal strategy $\varphi_{e_{i}}^{\mathrm{OPT}}$ that maximizes the LP reward $f_{e_{i}}^{\mathrm{LP}}$ within the liquidity of the approximated pool, cf.(22). Figure 8 illustrates how the liquidity state in each bucket changes while historical prices evolve, and $\bm{L}_{e_{i}}^{\Sigma}$ in the $2\tau+1$ buckets of $\beta$ contains, among other components, the LP-position in the form of $\bm{L}_{e_{i}}^{\mathrm{LP}}$ dictated by $\varphi_{e_{i}}^{\mathrm{OPT}}$.
@@
The distinguishing element of the present approach is that LP fees are calculated after local expansion of historical liquidity by the LP’s own liquidity $\bm{L}_{e_{i}}^{\mathrm{LP}}$ into $2\tau+1$ buckets centered on the reference index $M$. Thus, the fee share is taken relative to the augmented liquidity profile - historical plus the additional capital of the LP - see Fig. 10.
@@ 5 Modeling on Real Data
In this section, the methodology developed above is applied to real Uniswap V3 data on Ethereum. A uniform-liquidity strategy is used as a benchmark, while the proposed approach is evaluated on data from USDC/ETH pools with several fee levels, in stablecoin pools and in WBTC/ETH. The section ends with (i) a summary of the consolidated performance of all pools, (ii) a comparison with an external backtesting tool, and (iii) a high-level schematic of the application of the ML model for liquidity allocation decisions.
@@
Historical swap‑transaction data were retrieved using GraphQL queries to the Uniswap V3 subgraph.
@@ 5.1 Historical Liquidity Approximation and Optimal LP Strategy Search
For every epoch $e_{i}$ the pool’s liquidity shape is approximated by sweeping the Gaussian parameters $\mu_{e_{i}}\in[-3,3]$ and $\sigma_{e_{i}}\in[\sqrt{0.01},\sqrt{10}]$ until the reward of the modeled pool $f_{e_{i}}^{m\!\Sigma}(\mu_{e_{i}},\sigma_{e_{i}})$ deviates from the historical value $f_{e_{i}}^{\text{hist}}$ by no more than $5\%$. The average TVL is fixed at $\Sigma=80m\,$USDC, width of the bucket $d=10$ USDC, lower boundary $p_{a_{1}}=0$ and $N=650$. Applying the $\tau$-reset rule with $\tau=5$ partitions the in-sample period into $K=820$ epochs. The cumulative historical fee $\sum_{i=1}^{820}f_{e_{i}}^{\text{hist}}\approx 24.1m$ USDC is matched by the modeled fee $\sum_{i=1}^{820}f_{e_{i}}^{m\!\Sigma}\approx 23.8m$ USDC, that is, an approximation error of about $1.3\%$, well inside the $5\%$ tolerance. Figure 12 illustrates the calibration for the first epoch: the optimal parameters $\mu_{e_{1}}^{\ast}=-2.0$ and $\sigma_{e_{1}}^{\ast}=\sqrt{4.28}$ yield $f_{e_{1}}^{m\!\Sigma}=55\,348$ USDC versus historical 57 25957\,259 USDC (error $3.5\%$). The right-hand panel of Fig. 12 shows the resulting liquidity profile $\bm{L}_{e_{1}}^{\Sigma}$ and the active price range of epoch $e_{1}$ (in red).
@@
For epoch $e_{1}$ the optimizer produces the vector $\bm{W}_{e_{1}}^{\mathrm{LP}}$ shown in Fig. 11 (right), which aligns with the price path in the left panel and delivers a modeled LP fee of $f_{e_{1}}^{\mathrm{LP}}\bigl(\varphi_{e_{1}}^{\mathrm{OPT}}\bigr)=39\,162$ USDC.
@@ 5.2 ML Model Training and Features Engineering
Feature engineering is also intentionally minimal. All inputs are derived from only two raw numerical indicators, from which a compact set of features is generated to capture latent patterns in market dynamics. The approach is generic and can be adapted to any collection of numeric variables. The use of external market data for feature engineering is further justified by empirical studies showing that cryptocurrency markets react distinctively to news and sentiment compared to traditional asset classes [46], and exhibit unique, stress-sensitive volatility spillovers with conventional currencies [49]. Subsequent sections confirm that even such sparse feature sets and simple model configurations are sufficient for the goals of this work.
@@
The predictions of the integral model are produced by a weighted ensemble of three base learners: an MLP neural network, CatBoost, and an LSTM network. The remainder of this subsection details the architectures implemented in the experiments, including structural choices, theoretical rationale, hyperparameter settings, regularization techniques, and evaluation metrics. The use of an integrated ensemble model for prediction in this domain is supported by recent literature, which employs similar machine learning architectures to interpret and forecast complex, non-linear dynamics in DeFi and digital asset markets [48].
@@
[…]
@@
[…]
@@
The final prediction is obtained by a weighted sum of the three base models. The optimal weight vector $\bm{w}=\{w_{1},w_{2},w_{3}\}$ is found by minimizing the mean square error with regularization $L_{2}$ ($\alpha=0.001$) under the constraints $w_{j}>0$ and $\sum_{j=1}^{3}w_{j}=1$. Despite its architectural simplicity, the ensemble introduces methodological novelty by embedding structurally diverse models into a domain-specific backtesting framework for liquidity provision
@@ 5.3 Feature Engineering
Centralized exchange (CEX) data serve as the information set to characterize the market state in each allocation decision. Historical market data are sourced through the live_trading_indicators [^2] https://pypi.org/project/live-trading-indicators/ library, which offers unified access to multiple cryptocurrency exchanges. For the USDC/ETH pool, the study employs one-minute Open-High-Low-Close-Volume (OHLCV) bars for the USDC/USDT and BTC/USDT pairs on Binance, covering 1 March 2023 to 30 September 2024; the inclusion of BTC captures potential cross-asset comovement with ETH. For other pools, the same procedure is followed: features are generated for both the target pair and the BTC.
@@ 5.4 Application of ML Models to the OOT Sample
Figure 17: Accumulated LP fees depending on $\tau$
on the out-of-time period (September 2024) for USDC/ETH 0.3% pool.
@@
  • Gas expenditure. Deploying liquidity in a single range is charged 430 000430\,000 gas units, whereas burning liquidity costs 215 000215\,000 gas units. If the amount of LP liquidity in a given bucket remains unchanged after reallocation, the associated gas cost is ignored, in line with the convention of [6]. In this research, we do not consider the costs associated with token swaps necessary to convert into the desired asset for liquidity allocation. A constant gas price of $20\,\text{Gwei}$ is assumed.
@@
Throughout the OOT backtesting, all fees are not reinvested between epochs; this separation isolates the modeled reward stream from the evolving principal $W^{\mathrm{LP}}$. It is worth adding that impermanent loss in AMMs also can be interpreted as an AMM-specific manifestation of adverse selection known from traditional trading, where market makers systematically trade against better-informed (arbitrage) order flow.
@@
on the out-of-time period (September 2024) for USDC/ETH 0.3% pool
@@
Range Size ($d$) 10 USDC 10 USDC 1.23e-4 WBTC 1.57e-4 USDC
@@
5.5 Liquidity Management Example: "As-Is" and Modified $\tau$-reset Strategies
The primary focus of this research is on the quantitative assessment of LP rewards obtained through optimal liquidity placement based on historical data using ML models. However, as previously noted, this alone is insufficient for effective liquidity management as a whole. This is due, among other things, to a number of adverse factors discussed in Section 5.4, which are characteristic of both concentrated liquidity market makers (e.g., IL) and reset strategies in particular (e.g., adverse capital revaluation when the pool price crosses the lower bound of an LP’s active range - downside risk). We will refer to these as "natural" adverse factors. Furthermore, in order to obtain realistic out-of-time (OOT) backtesting results, it is necessary to account for the influence of other participants, rather than modeling all other LPs as a single static liquidity provider. Key examples of such influences include the following:
  • Just-in-Time (JIT) liquidity — a strategy in which other LPs add liquidity to the pool immediately prior to a swap execution to capture profit with minimal risk and then promptly remove it [29].
  • Maximal Extractable Value (MEV) — a practice used by block producers (validators or miners) to increase their revenue by selectively including, excluding, or reordering transactions when forming a block, includes frontrunning and so-called sandwich attacks [30]. It should be noted that JIT-strategy is part of MEV.
Although the examples listed above are full-fledged research topics in their own right, in the context of this work they serve merely as justification for the need to introduce additional assumptions in modeling. Ignoring interactions with other pool participants leads to an overestimation of strategy effectiveness during backtesting.
In this research, we take into account only the "natural" adverse factors. The left panel of Figure 19 shows the dynamics of active LP capital in USDC when applying the "As-Is" strategies developed in the USDC / ETH 0.05% pool, as a function of the parameter $\tau$. Modeling here is similar to the Section 5.4: the initial capital for all strategies is $W^{\mathrm{LP}}=1m\ \text{USDC}$, but with reinvestment of the rewards after each liquidity relocation (epoch). A clear inverse relationship is observed between capital preservation and the number of relocations, which is governed by the parameter $\tau$. Choosing $\tau$, the LP determines the capital concentration and thereby sets the price curve with specific levels of convexity, as described in [31].
Figure 19: The state of total liquidity positions in USDC depending on the $\tau$-parameter.
Application of "As-Is" $\tau$-reset strategies (integral models) on the USDC/ETH 0.05% pool
with width of the bucket $d=10$ USDC and $W^{\mathrm{LP}}=1m\ \text{USDC}$.
The right panel of Figure 19 also demonstrates an inverse relationship between the cumulative reinvested model LP rewards and the final level of LP capital in USDC at the end of the OOT period. Table 3 presents comparative results for the application of the optimal "As-Is" strategies and the buy and hold (B&H) strategy. Here, unlike the modeling of LP’s rewards alone, the buy and hold strategy serves as a natural benchmark. Additionally, we use classical performance metrics such as Max Drawdown and the Sharpe Ratio to evaluate and compare the effectiveness of the strategies.
Table 3: Analysis of strategies metrics depending on the $\tau$ parameter ("As-Is" integral models).
On the OOT period (September 2024) for the pool USDC/ETH 0.05 $\%$
$\tau$-parameter 0 11 22 55 $10$ 2020 $40$ B&H
Epochs (reallocations) 66466646* $641$ $241$ $55$ $16$ 44 11 th. USD
Costs $143.4$ $58.9$ $37.4$ $18.7$ $10.5$ $5.3$ $2.7$
Sum reinvest. fees $142.1$ $341.3$ $254.7$ $151.2$ $81.1$ $35.6$ $18.6$
LP capital $0.0$ $469.0$ $636.8$ $809.8$ $929.0$ $1008.4$ $1090.2$ $1035.5$
Compound annual returns 100.0-100.0% 100.0-100.0% 99.6-99.6% 92.0-92.0% 58.7-58.7% 10.510.5% 181.9181.9% 51.951.9%
Max Drawdown (MDD) 100.0-100.0% 53.3-53.3% 36.5-36.5% 19.9-19.9% 10.9-10.9% 10.3-10.3% 9.9-9.9% 16.2-16.2%
Sharpe Ratio (annual)** 13.5-13.5 7.7-7.7 3.1-3.1 $0.4$ $3.4$ $0.9$
* 16667 epochs without complete depletion of capital.
** with risk-free rate 0% and 365 crypto trading days per year.
The analysis conducted demonstrates the destructive impact of ill-considered liquidity management: excessive relocation frequency is a key driver of accelerated depletion of active LP capital—even when liquidity is placed optimally in theory and expected LP’s rewards are high, the strategy quickly loses its effectiveness.
Let us now consider in more detail the sensitivity of capital preservation to relocation frequency in the case of $\tau=5$, for which, according to the results of Table 2, the use of advanced ML models is more justified compared to a uniform strategy. To this end, we introduce several modifications of the $\tau$-reset strategy aimed at reducing sensitivity to frequent relocations while preserving the generalization ability of the ML model:
  • $\tau+\eta$ - reset strategy: in addition to the $2\tau+1$ LP-liquid buckets defined by the methodology in Section 2.2, $2\eta$ empty "side" ranges are introduced. Crossing into these ranges does not trigger a liquidity relocation. In this case, the set of LP-liquid buckets is $\beta_{L}^{\mathrm{LP}}=\{B_{M-\tau-\eta},\dots,B_{M+\tau+\eta}\}$, where the allocation weights $\alpha_{M-\tau-\eta+k}^{\mathrm{LP}}=\alpha_{M+\tau+\eta-k}^{\mathrm{LP}}=0$ for $k\in[0,\eta-1]$, and $M$ denotes the index of the reference bucket. The total coverage thus becomes $2(\tau+\eta)+1$ buckets. This symmetric strategy reduces excessive relocations during sharp price swings; however, capital remains subject to the "natural" adverse factors, although to a lesser extent.
  • $\tau+\eta^{\mathrm{up}}$ and $\tau+\eta^{\mathrm{down}}$ - reset strategies: in these asymmetric variants, the empty protective buckets are placed only in one direction: either toward the price increase or decrease, relative to the reference bucket. For instance, in the case of $\tau+\eta^{\mathrm{down}}$, we have $\beta_{L}^{\mathrm{LP}}=\{B_{M-\tau-\eta^{\mathrm{down}}},\dots,B_{M+\tau}\}$, with $\alpha_{M-\tau-\eta^{\mathrm{down}}+k}^{\mathrm{LP}}=0$ for $k\in[0,\eta^{\mathrm{down}}-1]$. The total coverage thus becomes in this case is $2\tau+\eta^{\mathrm{up/down}}+1$ buckets. When $\eta^{\mathrm{down}}>0$, the strategy reduces relocations during sharp price declines and utilizes capital more efficiently during price increases.
Table 4 presents a comparison of key metrics for the integral model using the $\tau+\eta^{\mathrm{down}}$-reset strategy with $\tau=5$, across different values of the $\eta^{\mathrm{down}}$ parameter:
Table 4: Analysis of strategies metrics depending on the $\eta^{down}$ parameter (integral model).
On the OOT period (September 2024) for the pool USDC/ETH 0.05 $\%$, $\tau=5$ and $W^{\mathrm{LP}}=1$ m USDC
$\eta^{down}$-parameter 0 55 $10$ $15$ 2020 $25$ $30$ $35$ $40$ B&H
Epochs (reallocations) $55$ $31$ 2121 $13$ 99 $11$ $11$ 44 44 th. USD
Positive realloc. 51.951.9% 66.766.7% 75.075.0% 83.383.3% 87.587.5% 90.090.0% 90.090.0% 100.0100.0% 100.0100.0%
Costs $18.7$ $10.7$ $7.2$ $4.5$ $3.2$ $3.8$ $3.8$ $1.5$ $1.5$
Sum reinvest. fees $151.2$ $119.7$ $95.4$ $72.7$ $86.6$ $85.6$ $65.6$ $35.9$ $35.9$
LP capital $809.8$ $868.5$ $917.0$ $989.4$ $1054.4$ $1009.7$ $999.1$ $1093.4$ $1093.4$ $1035.5$
Compound annual returns 92.0-92.0% 81.6-81.6% 64.7-64.7% 12.0-12.0% 88.988.9% 12.212.2% 1.1-1.1% 192.1192.1% 192.1192.1% 51.951.9%
Max Drawdown (MDD) 19.9-19.9% 17.9-17.9% 13.9-13.9% 16.3-16.3% 11.8-11.8% 13.5-13.5% 12.6-12.6% 12.9-12.9% 12.9-12.9% 16.2-16.2%
Sharpe Ratio (annual)* 7.7-7.7 3.8-3.8 2.3-2.3 0.2-0.2 $1.7$ $0.3$ $0.05$ $2.5$ $2.5$ $0.9$
* with risk-free rate 0% and 365 crypto trading days per year.
The results show that deliberately reducing sensitivity to sharp but temporary price drops leads to a decreased number of relocations (epochs), while the final capital at the end of the OOT period exhibits an upward trend. Moreover, as $\eta^{\mathrm{down}}$ increases, the number of relocations triggered by price increases surpasses those triggered by price declines. This further contributes to capital preservation, since relocations during price drops lead to the conversion of liquidity into a depreciating risky asset, thereby locking in losses—a manifestation of the second "natural" adverse factor. We observe that the application of the asymmetric strategy outperforms both the benchmark and the classical $\tau$-reset strategy, delivering more robust results when accounting for reallocation fees and "natural" adverse factors.
This approach maintains the predictive power of the ML model, unlike the naïve increase in $\tau$, which effectively causes the performance of the integral model to converge toward that of a uniform strategy (see Section 5.4). Figure 20 shows the dynamics of active LP capital under the $\tau+\eta^{\mathrm{down}}$ - reset strategy with $\eta^{\mathrm{down}}=20$.
Figure 20: The state of total liquidity positions in USDC. Application of $\tau+\eta^{\mathrm{down}}$ - reset strategies (integral model) on the USDC/ETH 0.05% pool with $\tau=5$, $\eta^{\mathrm{down}}=20$ and $W^{\mathrm{LP}}=1m\ \text{USDC}$.
As seen in the figure, the modified strategy utilizes capital efficiently: for a significant portion of time, the price remains within the LP-liquid buckets, while the empty buckets act as a protective buffer. This modification of the baseline strategy delivers better performance metrics compared to the buy and hold strategy, and unlike configurations with $\eta^{\mathrm{down}}=35$ or $40$, it remains more dynamic. It is important to note that the $\tau+\eta$, $\tau+\eta^{\mathrm{up}}$ and $\tau+\eta^{\mathrm{down}}$-reset strategies are merely extensions of the baseline $\tau$-reset strategy and thus do not require retraining of the ML models developed earlier. As a result of the analysis, the following conclusions can be formulated:
  • Reducing the frequency of relocations and prioritizing relocations during upward price movements are effective capital protection measures. However, these measures inherently favor static liquidity provision strategies, thereby diminishing the added value of using advanced ML models.
  • In the context of dynamic liquidity management, applying the methodology proposed in this research—alongside hedging mechanisms—appears to be a necessary condition for achieving sustainable returns and outperforming passive strategies. Otherwise, observed results may merely reflect coincidental market conditions and exhibit high volatility, similar to the behavior of the buy and hold strategy.
This extended problem formulation—constructing an integrated active liquidity placement strategy with hedging—will be the subject of our follow-up research.
5.6 Comparison with an Alternative Tool